# Dense 3D Displacement Estimation for Landslide Monitoring via Fusion of TLS Point Clouds and Embedded RGB Images

### v1

@@ Abstract
Landslide monitoring is essential for understanding geohazards and mitigating associated risks. However, existing point cloud-based methods typically rely on either geometric or radiometric information and often yield sparse or non-3D displacement estimates. In this paper, we propose a hierarchical partition-based coarse-to-fine approach that fuses 3D point clouds and co-registered RGB images to estimate dense 3D displacement vector fields. We construct patch-level matches using both 3D geometry and 2D image features. These matches are refined via geometric consistency checks, followed by rigid transformation estimation per match. Experimental results on two real-world landslide datasets demonstrate that our method produces 3D displacement estimates with high spatial coverage (79% and 97%) and high accuracy. Deviations in displacement magnitude with respect to external measurements (total station or GNSS observations) are 0.15 m and 0.25 m on the two datasets, respectively, and only 0.07 m and 0.20 m compared to manually derived references. These values are below the average scan resolutions (0.08 m and 0.30 m). Our method outperforms the state-of-the-art method F2S3 in spatial coverage while maintaining comparable accuracy. Our approach offers a practical and adaptable solution for TLS-based landslide monitoring and is extensible to other types of point clouds and monitoring tasks. Our example data and source code are publicly available at github.com/zhaoyiww/fusion4landslide.
@@ 1 Introduction
To address the aforementioned issues, we propose a deep learning-based, cross-modal approach to estimating dense 3D displacements for TLS-based landslide monitoring. Our key innovation lies in first performing hierarchical partition-based coarse (patch) matching, followed by fine matching between each matched patch pair. This process considers both point clouds and RGB images. To mitigate unreliable correspondences computed purely in Euclidean space, we instead adopt feature-based correspondences that use deep features extracted by pre-trained deep learning models with strong generalization (Poiesi and Boscaini, 2021), inspired by the strategy in F2S3 (Gojcic et al., 2021). Furthermore, to avoid the limitations of uniform partitioning, which can disrupt object boundaries by ignoring underlying structures, and single-level partitioning, which may sacrifice spatial coverage to reduce the number of noisy displacement estimates, we adopt an advanced hierarchical partitioning method. This method partitions input point clouds into three levels of partitions, taking geometric and (optionally) radiometric features into account. By optimizing the matching pipeline and fusing 3D point clouds and 2D RGB information, our approach achieves displacement estimation accuracy comparable to the current state-of-the-art method (F2S3) while improving spatial coverage.
@@ 2.1 Point cloud-based displacement estimation
Many methods have been proposed to estimate displacements from point clouds acquired at different epochs, To estimate displacements from point clouds, either from TLS scanners or from photogrammetric 3D reconstructions. Here we broadly classify these methods into five categories. Category I includes traditional methods such as C2C (Girardeau-Montaut et al., 2005), C2M (Cignoni et al., 1998), and original M3C2 (Lague et al., 2013), which estimate displacements by comparing spatial distances in Euclidean space, typically along surface normals. These methods operate via nearest neighbor (NN) searches or projection onto local reference surfaces, using strategies such as point-to-point, point-to-mesh, or averaged point-to-neighborhood comparisons. While effective in many cases, they are sensitive to variations in point density and surface roughness and often miss tangential displacements occurring along the surface. Several adaptations James et al. (2017); Zahs et al. (2022); Yang and Schwieger (2023) aim to improve correspondence quality or enhance the selection of informative directional estimates. Multi-directional M3C2 (Williams et al., 2021) improves flexibility by considering multiple measurement directions. However, without explicit point-wise correspondence tracking, these methods estimate displacement only along predefined direction(s) rather than recovering full 3D displacement vectors.
Category II entails “Piecewise ICP” and related alternatives (Teza et al., 2007; Friedli and Wieser, 2016; Wujanz et al., 2018), which uniformly partition point clouds into tiles and estimate per-tile displacements as the norms of translation vectors obtained through ICP algorithm (Besl and McKay, 1992; Chen and Medioni, 1992; Bergevin et al., 1996) or some of its variants. In this case, the per-tile correspondences are also established directly in the Euclidean space. Despite their effectiveness in certain scenarios, this class of methods produces yields biased estimates in instances where per-tile rigid body motion is violated or points within the tile are inadequate for accurate estimation.
@@
Category V encompasses methods developed for point clouds generated through photogrammetric 3D reconstruction rather than TLS. For instance, Wagner (2016) employs a total station equipped with a telescope camera to obtain 3D observations and RGB images simultaneously. Matching is then conducted on RGB images using cross-correlation methods. Lucks et al. (2024) applies SIFT (Lowe, 2004) to UAV images to identify 2D matches, which are then reconstructed into 3D displacements using photogrammetric techniques. Similarly, Ioli et al. (2023, 2024) perform dense 3D reconstruction from UAV data for glacier monitoring. They utilize deep-learning-based image matching algorithms for matching, leveraging deep-learning-based image matching to yield denser displacement vectors than earlier approaches. While these methods are relatively easy to implement, their accuracy heavily depends on reconstruction quality and they are often sensitive to large illumination variations and seasonal changes.
@@
Instead of explicitly detecting and describing keypoints and then matching them, detector-free methods integrate feature matching directly into the learning process. These methods typically employ a coarse-to-fine step to implement matching (Li et al., 2020; Zhou et al., 2021). More recently, LoFTR (Sun et al., 2021) and its variant Efficient LoFTR (Wang et al., 2024), introduce transformer-based architectures to improve the accuracy of 2D image matching. Inspired by the coarse-to-fine matching mechanism used for 2D image matching, CoFiNet (Yu et al., 2021) adopt a similar mechanism for the 3D point matching task. Building on this framework, GeoTransformer (Qin et al., 2022) incorporate transformer networks to further improve registration accuracy, particularly in challenging low-overlap scenarios.
Our proposed method is detector-free and introduces two key innovations over existing detector-free approaches: (i) Existing methods rely on fully convolutional networks that process entire point clouds, which limits their applicability to large-scale datasets. In contrast, our method leverages a local patch-based network (Poiesi and Boscaini, 2021) with strong generalizability. This design is especially suited for landslide monitoring, where point clouds may contain tens of millions of points and labeled training data are scarce. (ii) Existing methods often employ uniform subsampling to generate hierarchical keypoints or patches (Thomas et al., 2019). Instead, we adopt a hierarchical, feature-based partitioning strategy that aligns better with object boundaries, enabling more effective motion pattern estimation.
@@ 3 Methodology
Figure 1: Method overview. The proposed pipeline consists of five steps: (1) Point cloud tiles and the corresponding image tiles are generated. (2) Each paired source-target point cloud tile is partitioned into smaller patches. (3) Coarse matching establishes patch-wise correspondences / matches based on both 3D point cloud and 2D RGB information. (4) A refinement module evaluates match quality and discards low-quality patch matches based on predefined criteria. (5) For each pair of patch matches, a rigid transformation is estimated and applied to all points within the source patch. The transformed source points and associated displacement vectors are then aggregated to construct the final 3D DVF, optionally producing point-wise matches.
@@
To estimate dense 3D DVF for landslide monitoring, we design a fusion-driven approach that jointly leverages TLS point clouds and co-registered RGB images. The RGB images, captured by built-in camera(s), are co-registered with the point cloud at the respective epoch (see details in Sec. 4.1). The proposed method consists of five key steps: tiling large point clouds and the corresponding RGB images (cf. Sec. 3.1); partitioning each point cloud tile into smaller patches (cf. Sec. 3.2); generating candidate coarse matches between patches (cf. Sec. 3.3); refining these coarse matches (cf. Sec. 3.4), and performing fine matching (cf. Sec. 3.5). An overview of the proposed method is illustrated in Fig. 1. While designed for two epochs, the method is easily generalizable to multi-epoch scenarios by sequentially processing pairs of epochs.
@@ 3.1 Point cloud and image tiling
Due to the large size of point clouds (e.g., tens of millions of points per epoch) representing an entire landslide, we divide the point clouds into smaller spatial tiles to improve computational efficiency and reduce memory usage. Following Gojcic et al. (2021), we first project the source and target point clouds to 2D along the coordinate axis that yields the largest bounding box in the projection. This bounding box is then recursively subdivided along the longer edge until each tile contains fewer than a predefined number of points (e.g., 1 million). Each tile pair, consisting of a source and a target tile that spatially overlap the same 3D region, is processed independently of the other tile pairs in the subsequent steps of our method. To accommodate displacements that may span tile boundaries, adjacent tiles are defined with an overlap margin equal to the maximum expected displacement (e.g., 10 m, set by a manually adjustable threshold). To ensure computational efficiency, the RGB images are divided into smaller overlapping tiles before being processed by the image matching algorithm.
3.2 Hierarchical point cloud partition
@@
The local rigidity assumption is commonly applied in many applications (Sorkine and Alexa, 2007; Yang et al., 2014; Dewan et al., 2016). We argue that this assumption also holds for landslide cases, as small objects are more likely to move as rigid bodies. For example, individual boulders are expected to move as a single unit from one location to another, rather than splitting into separate pieces. To leverage this property, we propose partitioning the tiled point clouds into small patches, where patches correspond to those small objects or small regions undergoing locally consistent motion. Selecting an appropriate patch size is critical for accurate displacement estimation. If the patch size is too large, it may encompass multiple objects undergoing non-rigid motion, which breaks the local rigidity assumption. Conversely, patches that are too small are prone to containing mainly noise.
Figure 2: Hierarchical partitioning of the point cloud. Starting from the input point cloud colorized by RGB values, the method generates three levels of partitions that capture structural boundaries at multiple scales. Level i consists of the smallest and most numerous patches, while level iii contains the largest and fewest patches.
@@
To capture multi-scale structural variations in the scene, we adopt the hierarchical partitioning strategy proposed by Robert et al. (2023), which decomposes the point cloud into patches at three levels of granularity (cf. Fig. 2) using pre-computed handcrafted features. We incorporate both geometric (linearity, planarity, and curvature) and radiometric (grayscaled RGB) features to guide this partitioning. Each partition level is used independently to guide patch-wise matching in the subsequent coarse-to-fine pipeline. At a later stage (cf. Sec. 3.5), we integrate the displacement estimates from all three levels to maintain high accuracy and high spatial coverage. Additionally, to mitigate the impact of noise, patches containing fewer than ten points are discarded for matching.
@@ 3.3 Candidate generation for coarse matching
For coarse matching, also known as patch matching, we define a match between a source and target patch as the pair with the highest similarity based on their geometric and radiometric attributes. A match may still be declared even if the patches (i) differ in their shape or (ii) undergo local orientation changes relative to their neighbors, as illustrated in Fig. 3. However, this matching procedure is impacted by the fact that the partitioning method rarely generates identical patches for the same regions across two epochs, even when the regions remain stable. To address this challenge and ensure accurate matching between patches from different epochs, we develop a processing scheme that is robust with respect to (w.r.t.) the exact patch shapes. This scheme contains patch matching derived from 3D source data (cf. Sec. 3.3.1) and patch matching derived from 2D source data (cf. Sec. 3.3.2). The final set of initial coarse matches combines both results, ensuring a balanced use of different types of information.
@@ 3.3.1 Patch matching from 3D source
Patch matching from the 3D source (i.e., point cloud) relies on 3D point feature matching. To ensure feature discriminativeness, we first downsample the point cloud tiles to achieve a uniform density. This prevents feature descriptors from over-focusing on densely populated regions while neglecting sparse areas. Unlike methods such as F2S3 (Gojcic et al., 2021)), which use a fixed voxel size for downsampling and feature extraction across all tiles, we adjust the voxel size adaptively based on the local density (i.e., the average scan resolution) of each tile. Next, we use a pre-trained local patch-based 3D neural network (Poiesi and Boscaini, 2021) to extract 3D point features from each point in the patch. Since the objective is to establish patch-wise matches, we then focus on aggregating the features of all points within each patch.
Figure 4: Coarse matching based on patch feature similarity. (i) Network architecture for constructing patch features using point features extracted from a pre-trained model within each patch. (ii) Mutual NN matching between patch features from the two epochs for generating candidate patch matches.
To achieve this, we design a lightweight network (cf. Fig. 4) and train it on publicly available indoor datasets (Zeng et al., 2017), eliminating the need for effort in annotating landslide training data. After the pre-trained model assigns features to the points within a patch, our network begins by using self-attention layers (Vaswani, 2017) to adaptively aggregate these features. The aggregated features then go through MLP layers to form the final patch features. Patches from the source and target point clouds are processed identically to construct patch features. During training, we apply the circle loss (Sun et al., 2020), a variant of the triplet loss (Choy et al., 2019) commonly used in 3D feature matching (Huang et al., 2021; Qin et al., 2022), to optimize the patch feature aggregation. In the testing phase, i.e., when estimating displacements for the real-world landslides, patch features are extracted using our pre-trained network. Patch-to-patch matches from the 3D source are then established through mutual NN matching based on their features.

### v2

@@ Abstract
Landslide monitoring is essential for understanding geohazards and mitigating associated risks. Existing point cloud-based methods, however, typically rely on either geometric or radiometric information and often yield sparse or non-3D displacement estimates. In this paper, we propose a hierarchical partitioning-based coarse-to-fine approach that integrates 3D point clouds and co-registered RGB images to estimate dense 3D displacement vector fields. Patch-level matches are constructed using both 3D geometry and 2D image features, refined via geometric consistency checks, and followed by rigid transformation estimation per match. Experimental results on two real-world landslide datasets demonstrate that the proposed method produces 3D displacement estimates with high spatial coverage (79% and 97%) and accuracy. Deviations in displacement magnitude with respect to external measurements (total station or GNSS observations) are 0.15 m and 0.25 m on the two datasets, respectively, and only 0.07 m and 0.20 m compared to manually derived references, all below the mean scan resolutions (0.08 m and 0.30 m). Compared with the state-of-the-art method F2S3, the proposed approach improves spatial coverage while maintaining comparable accuracy. The proposed approach offers a practical and adaptable solution for TLS-based landslide monitoring and is extensible to other types of point clouds and monitoring tasks.
@@ 1 Introduction
To address the aforementioned issues, we propose a deep learning-based, cross-modal approach for estimating dense 3D displacements in TLS-based landslide monitoring. The key innovation lies in hierarchical partitioning-based two-stage matching strategy, which performs coarse patch matching followed by fine matching within each matched patch pair. This process jointly considers both 3D point cloud geometry and 2D RGB image information. Specifically, to mitigate unreliable correspondences computed purely in Euclidean space, we instead adopt feature-based correspondences that use deep features extracted by pretrained deep learning models with strong generalization (Poiesi and Boscaini, 2021), inspired by the strategy in F2S3 (Gojcic et al., 2021). To overcome the limitations of uniform or single-level partitioning, which can disrupt object boundaries or sacrifice spatial coverage, we employ a three-level hierarchical partitioning scheme that incorporates geometric and (optionally) radiometric features. By optimizing the matching pipeline and fusing 3D point clouds and 2D RGB information, our approach achieves displacement estimation accuracy comparable to the current state-of-the-art method (F2S3) while significantly improving spatial coverage.
@@ 2.1 Point cloud-based displacement estimation
Many methods have been proposed to estimate displacements from point clouds acquired at different epochs, either from TLS scanners or from photogrammetric 3D reconstructions. Here we broadly classify these methods into five categories. Category I includes traditional methods such as C2C (Girardeau-Montaut et al., 2005), C2M (Cignoni et al., 1998), and original M3C2 (Lague et al., 2013), which estimate displacements by comparing spatial distances in Euclidean space, typically along surface normals. These methods operate via nearest neighbor (NN) search or projection onto local reference surfaces, using strategies such as point-to-point, point-to-mesh, or averaged point-to-neighborhood comparisons. While effective in many cases, they are sensitive to variations in point density and surface roughness and often miss tangential displacements occurring along the surface. Several adaptations James et al. (2017); Zahs et al. (2022); Yang and Schwieger (2023) aim to improve correspondence quality or enhance the selection of informative directional estimates. Multi-directional M3C2 (Williams et al., 2021) improves flexibility by considering multiple measurement directions. However, without explicit point-wise correspondence tracking, these methods estimate displacement only along predefined direction(s) rather than recovering full 3D displacement vectors.
Category II entails “Piecewise ICP” and related alternatives (Teza et al., 2007; Friedli and Wieser, 2016; Wujanz et al., 2018), which uniformly partition point clouds into tiles and estimate per-tile displacements as the norms of translation vectors obtained through ICP algorithm (Besl and McKay, 1992; Chen and Medioni, 1992; Bergevin et al., 1996) or some of its variants. In this case, the per-tile correspondences are also established directly in the Euclidean space. Despite their effectiveness in certain scenarios, this class of methods yields biased estimates in instances where per-tile rigid body motion is violated or points within the tile are inadequate for accurate estimation.
@@
Category V encompasses methods developed for point clouds generated through photogrammetric 3D reconstruction rather than TLS. For instance, Wagner (2016) employs a total station equipped with a telescope camera to obtain 3D observations and RGB images simultaneously. Matching is then conducted on RGB images using cross-correlation methods. Lucks et al. (2024) applies SIFT (Lowe, 2004) to UAV images to identify 2D matches, which are then reconstructed into 3D displacements using photogrammetric techniques. Similarly, Ioli et al. (2023, 2024) perform dense 3D reconstruction from UAV data for glacier monitoring. They utilize deep learning-based image matching algorithms for matching, leveraging deep learning-based image matching to yield denser displacement vectors than earlier approaches. While these methods are relatively easy to implement, their accuracy heavily depends on reconstruction quality and they are often sensitive to large illumination variations and seasonal changes.
@@
Instead of explicitly detecting and describing keypoints and then matching them, detector-free methods integrate feature matching directly into the learning process. These methods typically employ a coarse-to-fine step to implement matching (Li et al., 2020; Zhou et al., 2021). More recently, LoFTR (Sun et al., 2021) and its variant Efficient LoFTR (Wang et al., 2024), introduce transformer-based architectures to improve the accuracy of 2D image matching. Inspired by the coarse-to-fine matching mechanism used for 2D image matching, CoFiNet (Yu et al., 2021) adopts a similar mechanism for the 3D point matching task. Building on this framework, GeoTransformer (Qin et al., 2022) incorporates transformer networks to further improve registration accuracy, particularly in challenging low-overlap scenarios.
Our proposed method is detector-free and introduces two key innovations over existing detector-free approaches: (i) Existing methods rely on fully convolutional networks that process entire point clouds, which limits their applicability to large-scale datasets. In contrast, our method leverages a local patch-based network (Poiesi and Boscaini, 2021) with a strong generalizability. This design is especially suited for landslide monitoring, where point clouds may contain tens of millions of points and labeled training data are scarce. (ii) Existing methods often employ uniform subsampling to generate hierarchical keypoints or patches (Thomas et al., 2019). Instead, we adopt a hierarchical, feature-based partitioning strategy that aligns better with object boundaries, enabling more effective motion pattern estimation.
@@ 3 Methodology
Figure 1: Method overview. The proposed pipeline consists of five steps: (1) Point cloud tiles and the corresponding image tiles are generated. (2) Each paired source–target point cloud tile is partitioned into smaller patches. (3) Coarse matching establishes patch-wise correspondences / matches for each partition level based on both 3D point cloud and 2D RGB information. (4) A refinement module evaluates match quality and discards low-quality patch matches based on predefined criteria. (5) For each pair of patch matches, a rigid transformation is estimated and applied to all points within the source patch. The transformed source points and associated displacement vectors are then aggregated to construct 3D DVF, optionally producing point-to-point (P2P) matches. The final 3D DVF and P2P matches are integrated from different partition levels according to an integration step (cf. Section 3.5).
@@
To estimate dense 3D DVF for landslide monitoring, we design a fusion-driven approach that jointly leverages TLS point clouds and co-registered RGB images. The RGB images, captured by built-in camera(s), are co-registered with the point cloud at the respective epoch (see details in Section 4.1). The proposed method consists of five key steps: tiling large point clouds and the corresponding RGB images (cf. Section 3.1); partitioning each point cloud tile into smaller patches (cf. Section 3.2); generating candidate coarse matches between patches (cf. Section 3.3); refining these coarse matches (cf. Section 3.4), and performing fine matching (cf. Section 3.5). An overview of the proposed method is illustrated in Figure 1. While designed for two epochs, the method is easily generalizable to multi-epoch scenarios by sequentially processing pairs of epochs.
@@ 3.1 Point cloud and image tiling
Due to the large size of point clouds (e.g., tens of millions of points per epoch) representing an entire landslide, we divide the point clouds into smaller spatial tiles to improve computational efficiency and reduce memory usage. Following Gojcic et al. (2021), we first project the source and target point clouds to 2D along the coordinate axis that yields the largest bounding box in the projection. This bounding box is then recursively subdivided along the longer edge until each tile contains fewer than a predefined number of points, e.g., 1 million [^1] The maximum number of points per tile was set to avoid GPU memory overflow during processing. Tiles exceeding this limit are recursively subdivided. Empirically, a limit of 1 million points per tile works well on a single GeForce RTX 3090 Ti (24 GB) with an AMD Ryzen 7 5800X (8 cores). This parameter can be adjusted according to available memory capacity.. Each tile pair, consisting of a source and a target tile that spatially overlap the same 3D region, is processed independently of the other tile pairs in the subsequent steps of our method. To accommodate displacements that may span tile boundaries, adjacent tiles are defined with an overlap margin equal to the maximum expected displacement (e.g., 10 m, set by a manually adjustable threshold). To ensure computational efficiency, the RGB images are divided into smaller overlapping tiles before being processed by the image matching algorithm.
3.2 Hierarchical point cloud partitioning
@@
The local rigidity assumption is commonly applied in many applications (Sorkine and Alexa, 2007; Yang et al., 2014; Dewan et al., 2016). We argue that this assumption also holds for landslide cases, as small objects are more likely to move as rigid bodies. For example, individual boulders are expected to move as a single unit from one location to another, rather than splitting into separate pieces. To exploit this property, we propose partitioning the tiled point clouds into small patches, where patches correspond to those small objects or small regions undergoing locally consistent motion. Selecting an appropriate patch size is critical for accurate displacement estimation. Oversized patches may span multiple objects with non-rigid motions, violating the local rigidity assumption, whereas overly small patches tend to be dominated by noise.
Figure 2: Hierarchical partitioning of the point cloud. Starting from the RGB-colored input, the method generates three levels of partitions that capture structural boundaries at multiple scales. The levels are generated independently using different regularization strengths: Level i (or level 1) preserves small, geometrically homogeneous parts; Level $i$ reveals larger structural components; and Level iii highlights entire objects or coherent regions, providing global contextual information (Robert et al., 2023).
To capture multi-scale structural variations in the scene, we adopt the hierarchical partitioning strategy, which decomposes the point cloud into three levels of granularity (cf. Figure 2). The method relies on pre-computed geometric (linearity, planarity, and curvature) and radiometric (grayscaled RGB) features to ensure that the partitions align with both structural and appearance boundaries. Formally, for each tiled point cloud, the partitioning process applies a sequence of graph-based optimization problems with increasing regularization parameters that control the coarseness of the resulting patches (Robert et al., 2023). This yields multi-scale subsets defined as:
$\mathcal{P}^{(l)}=\{P^{(l)}_{k}\mid k=1,\dots,N_{l}\},\quad\mathcal{Q}^{(l)}=\{Q^{(l)}_{k}\mid k=1,\dots,N_{l}\},\quad l\in\{1,2,3\},$ (2)
where $P^{(l)}_{k}$ and $Q^{(l)}_{k}$ denote the $k$-th patches of the tiled source and target point clouds, $\mathcal{P}^{(l)}$ and $\mathcal{Q}^{(l)}$, respectively, at hierarchy level $l$.
@@
Each partition level is employed independently to guide patch-wise matching in the subsequent coarse-to-fine pipeline. At a later stage (cf. Section 3.5), we integrate the displacement estimates from all three levels to maintain high accuracy and high spatial coverage. To reduce noise, patches containing fewer than ten points are excluded from the matching process.
@@ 3.3 Candidate generation for coarse matching
For coarse matching, also known as patch matching, we define a match between a source and target patch as the pair with the highest similarity based on their geometric and radiometric attributes. A match may still be declared even if the patches (i) differ in their shape or (ii) undergo local orientation changes relative to their neighbors, as illustrated in Figure 3. However, this matching procedure is impacted by the fact that the partitioning method rarely generates identical patches for the same regions across two epochs, even when the regions remain stable. To address this challenge and ensure accurate matching between patches from different epochs, we develop a processing scheme that is robust with respect to (w.r.t.) the exact patch shapes. This scheme contains patch matching derived from 3D source data (cf. Section 3.3.1) and patch matching derived from 2D source data (cf. Section 3.3.2). The final set of initial coarse matches combines both results, ensuring a balanced use of different types of information.
@@ 3.3.1 Patch matching from 3D source
Patch matching from the 3D source (i.e., point cloud) relies on 3D point feature matching. To ensure feature discriminativeness, we first downsample the point cloud tiles to achieve a uniform density. This prevents feature descriptors from over-focusing on densely populated regions while neglecting sparse areas. Unlike methods such as F2S3 (Gojcic et al., 2021), which use a fixed voxel size for downsampling and feature extraction across all tiles, we adjust the voxel size adaptively based on the local density (i.e., the mean scan resolution) of each tile. Next, we use a pretrained local patch-based 3D neural network (Poiesi and Boscaini, 2021) to extract 3D point features from each point in the patch. Since the objective is to establish patch-wise matches, we then focus on aggregating the features of all points within each patch.
Figure 4: Coarse matching based on patch feature similarity. (i) Network architecture for constructing patch features using point features extracted from a pretrained model within each patch. (ii) Mutual NN matching between patch features from the two epochs for generating candidate patch matches.
To achieve this, we design a neural network (cf. Figure 4) that incorporates self-attention and MLP layers, and train it on publicly available indoor datasets (Zeng et al., 2017), eliminating the need for effort in annotating landslide training data. After the pretrained model assigns features to the points within a patch, our network begins by using self-attention layers (Vaswani, 2017) to adaptively aggregate these features. The aggregated features then go through MLP layers to form the final patch features. Patches from the source and target point clouds are processed identically to construct patch features. During training, we apply the circle loss (Sun et al., 2020), a variant of the triplet loss (Choy et al., 2019) commonly used in 3D feature matching (Huang et al., 2021; Qin et al., 2022), to optimize the patch feature aggregation. In the testing phase, i.e., when estimating displacements for the real-world landslides, patch features are extracted using our pretrained network. Patch-to-patch matches from the 3D source are then established through mutual NN matching based on their features.